OpenStudy (anonymous):

7. Given 3x - y = 15, find y when x = -4 A. 27 B. 3 C. -3 D. -27 8. Given 3x - y = 15, find y when x = -2 A. -21 B. -9 C. 9 D. 21 9. Given 3x - y = 15, find y when x = 2 A. -21 B. -9 C. 9 D. 21 10. Given 3x - y = 15, find y when x = 4 A. -27 B. -3 C. 3 D. 27

OpenStudy (anonymous):

Plug in the given value and solve. I'll do the first one. Given 3x - y = 15, find y when x = -4 3(-4) - y = 15 -12 - y = 15 Add 12. -y = 27 Divide by -1. y = -27 Try the rest and tell me what you get.